Calculate the equivalent capacitance between points A and B for the circuit shown. All the capacitors are 5 μF each. Give your answer in μF.

B —
Answer #1

two capacitors in bottom branch are in series, they give us

1/C = 1/5 + 1/5 = 2.5 uF

this is in parallel with middle capacitor in upper branch, we get

C = 2.5 + 5 = 7.5 uF

this is in series with left and right capacitors in upper branch

so,

Equivalent capacitance

1/C = 1/7.5 + 1/5 + 1/5

C = 1.875 uF